OpenStudy (anonymous):

solve for angle B. b=9, a=6, c=4

OpenStudy (lucaz):

any diagram?

OpenStudy (campbell_st):

the formula for the law of consines in this question is \[\cos(B) = \frac{a^2 +c^2 - b^2}{2ac}\] sorry about the mistake before

OpenStudy (anonymous):

i got \[Cos(B)=a ^{2}+7^{2}+9^{2}/2(a)(9)\]

OpenStudy (campbell_st):

it should be \[\cos(B) = \frac{6^2 + 4^2 - 9^2}{2 \times 6 \times4}\] and looking at the diagram, it will be the largest angle, since its opposite the longest side

OpenStudy (anonymous):

a is not 6, you don't know a from what i'm seeing

OpenStudy (campbell_st):

so when you get the answer... the next step is to find the angle... so it would be \[B = \cos^{-1}(answer)..... or.......B = \arccos(answer)\] depending on what you use to find an angle

OpenStudy (lucaz):

opposite side of angle A is 6, a=6

OpenStudy (anonymous):

i was looking at the wrong problem, yes a is 6 but i'm still confused haha i don't get this at all

OpenStudy (lucaz):

campbell showed you the law of cosine, you just have to replace the variables

OpenStudy (campbell_st):

well the law of cosines is a standard formula... the question is just about identifying the information and substituting it into the correct place,

OpenStudy (anonymous):

i got -29/48=cos(B)

OpenStudy (campbell_st):

you'll need to use brackets my method for input into a calculator is \[(6^2 + 4^2 - 9^2) \div (2 \times 4 \times 6)\] then you'll get a negative fraction answer...

OpenStudy (campbell_st):

great so on your calculator press shift or 2nd then cos then ANS close the brackets press equals and you should get the answer.

OpenStudy (campbell_st):

then ANS or answer keay just means you don't need to input the fraction again.

OpenStudy (anonymous):

i don't think i'm right.. i got 0.82298

OpenStudy (campbell_st):

not quite so what sort of calculator do you have...?

OpenStudy (anonymous):

i forgot my calculator at school so at the moment i'm using mathway.com

OpenStudy (campbell_st):

ok... hold on this online calculator will work. http://web2.0calc.com/ press 2nd then acos then +- 29/48 close brackets... and lastly the equals

OpenStudy (anonymous):

127.168899655999

OpenStudy (anonymous):

then to round that to the nearest 10th it would be 127.2 right?

OpenStudy (campbell_st):

yep thats great... well done

OpenStudy (anonymous):

thank you so much! these take home tests are so hard for me

OpenStudy (campbell_st):

well remember if it take home... work with the formulas next to you... good luck